On​ Monday, Harold picked up six donuts and five large coffees for the office staff. He paid $ 6.86 .  On​ Tuesday, Melinda picked up two donuts and four large coffees for the office staff. She paid $ 4.06 .  What is the cost of one​ donut? What is the cost of one large​ coffee?
Answer by stanbon(75887)

On​ Monday, Harold picked up six donuts and five large coffees for the office staff. He paid $ 6.86 .  On​ Tuesday, Melinda picked up two donuts and four large coffees for the office staff. She paid $ 4.06 .  What is the cost of one​ doughnut? What is the cost of one large​ coffee?
-------------
Equations:
6d + 5c = 6.86
2d + 4c = 4.06
---------------------
Modify to solve for "d"::
24d + 20c = 4*6.86
10d + 20c = 5*4.06
-------------------------
Subtract and solve for "d"::
14d = 7.14
d = $0.51 (price of one doughnut)
Solve for "c"::
2d + 4c = 4.06
102 + 4c = 406
4c = 304
c = 76 cents = $0.76 (price of one coffee)
--------------------
Cheers,
Stan H.
------------------